State expression x+1/x(x+5) as a partial fractions and determine the values of constans.
Answer by algebrapro18(249)

x+1/x(x+5) = A/x + B/(x+5)
x+1 = A(x+5)+B(x)
let x = 0
1=5A
A= 1/5
let x = -5
-4=-5B
4/5 = B
x+1/x(x+5) = 1/5/x + 4/5/(x+5) = 1/5x + 4/5(x+5)